how much of an 80% nitric acid solution must be added to a 125gram solution of 10% nitric acid to produce a 50% nitric acid solution?
Answer by stanbon(57984)

how much of an 80% nitric acid solution must be added to a 125gram solution of 10% nitric acid to produce a 50% nitric acid solution?
-----------------------
Equation:
active + active = active
0.10*125 + 0.80x = 0.50(125+x)
---
Multiply thru by 100 to get:
10*125 + 80x = 50*125 + 50x
30x = 40*125
x = 166 2/3 grams (amt. of 80% solution needed in the mixture)
==============================
Cheers,
Stan H.